OpenStudy (anonymous):

A certain forest covers an area of 4200km2. Suppose that each year this area decreases by 5.5%. What will the area be after 12 years?

OpenStudy (campbell_st):

are you working on sequences and series...?

OpenStudy (anonymous):

idk @campbell_st

OpenStudy (campbell_st):

ok... makes it hard to explain, if you don't know. do you know how to use the compound interest formula..?

OpenStudy (anonymous):

yes but i need it to decrease and idk how to do that one @campbell_st

OpenStudy (campbell_st):

ok... so the compound interest formula can be converted to a depreciation formula by changing the sign inside the brackets from + to -

OpenStudy (campbell_st):

the formula I know is \[A = P(1 + \frac{r}{100})^n\] you can use \[A = P(1 - \frac{r}{100})^n\]

OpenStudy (campbell_st):

r = 5.5 P = 4200 and n = 12 hope it helps

OpenStudy (anonymous):

thank you i will try it @campbell_st

OpenStudy (anonymous):

i cant get it right @campbell_st

OpenStudy (campbell_st):

why, what do you get as an answer..?